Write the equation of the line that passes through point (6,4) with a slope of 2

Thanks

Found 2 solutions by stanbon, venugopalramana:
Answer by stanbon(75887)

Write the equation of the line that passes through point (6,4)
with a slope of 2
y=mx+b
You need to find m and b.
You are told that m=2
You substitute for x,y, and m to find b, as follows:
4=2(6)+b
b=-8
EQUATION:
y=2x-8
Cheers,
Stan H.

Answer by venugopalramana(3286)

Write the equation of the line that passes through point (6,4) =(X1,Y1)with a slope of 2 =M
Y-Y1=M(X-X1)
Y-4=2(X-6)
Y=2X-8